On average, 8 eligble aaulits) ut of every 4 gives blood The standard deviation is 60, so in most samples of four eligiblo adulits, the number of adults who donate blood would dfmes trom the avorage umbor by no more than (Type megers or decimals rounded to the nearest hundredth as needed) Enter your answer in the edit fields and then click Check Answer
4.2.29 Question Help * Twenty-two percent of people in a city eligible to donate biood actually do You randomly select four eligble blood donors and ask them i they donate blood Complote parts (a) through (d) below (c) Find the 83 to the nearest hundredth as needed ) On avorage 88 ligble adults) out of every 4 gives blood The standard devation e6, so in most samples of four eligble adults,the nuba aduts th domato tlood would diffor from the average number by no more than |0 t adults who donate blood would (Type intogers or decimals rounded to the nearest hundredth as needed.) Enter your answer in the edit fields and then click Check Answer All parts showing Clear All
Need help with part D
Answer #1

n = 4, p = 0.22

Mean n*p-4 * 0.22-0.88

Standard deviation = Vn * p * (1-p) = V4 * 0.22 * 0.78 0.828

On average 0.88 eligible adult(s) out of every 4 gives blood. The standard deviation is 0.828, so in most samples of four eligible adults, the number of adults who donate blood would differ from the average by no more than 0.